# Nyström Kernel Stein Discrepancy

### v2

@@ 4.2 Nyström bootstrap testing
$w_{i}=\mathds{1}_{(U_{i}>0.5)}w_{i-1}-\mathds{1}_{(U_{i}\leq 0.5)}w_{i-1},$ where $U_{i}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\operatorname{Unif}(0,1),$ (14)
that is, $w_{i}$ changes sign with probability $0.5$. The test procedure is as follows.
@@
with $\mathbf{w}=\left(w_{i}\right)_{i=1}^{n}\in\mathbb{R}^{n}$ collecting the $w_{i}$-s ($i\in[n]$) defined in (14); $\mathbf{K}_{h_{p},n,m}$ $(=\mathbf{K}_{h_{p},m,n}^{\mkern-1.5mu\mathsf{T}})$ and $\mathbf{K}_{h_{p},m,m}$ are defined as in Lemma 1. The approximation is based on the fact [Williams and Seeger, 2001] that $\mathbf{K}_{h_{p},n,m}\mathbf{K}_{h_{p},m,m}^{-}\mathbf{K}_{h_{p},m,n}$ is a low-rank approximation of $\mathbf{K}_{h_{p},n,n}$, that is, $\mathbf{K}_{h_{p},n,m}\mathbf{K}_{h_{p},m,m}^{-}\mathbf{K}_{h_{p},m,n}\approx\mathbf{K}_{h_{p},n,n}$. Hence, our proposed procedure (15) approximates (13) but reduces the cost from $\mathcal{O}\left(n^{2}\right)$ to $\mathcal{O}\left(nm+m^{3}\right)$ if one computes from left to right (also refer to Remark 1); in the case of $m=o\left(n^{2/3}\right)$ this guarantees an asymptotic speedup. We obtain a total cost of $\mathcal{O}\left(D\left(nm+m^{3}\right)\right)$ for obtaining the wild bootstrap samples. This acceleration allows KSD-based goodness-of-fit tests to be applied on large data sets.
@@ 4.3 Guarantees
Theorem 1.
Assume the setting of Lemma 1, $C_{\mathbb{Q},h_{p}}\neq 0$, $m\geq 4$ Nyström samples, and bounded Stein feature map ($\sup_{\mathbf{x}\in\mathbb{R}^{d}}\left\|{h_{p}\left(\cdot,{\mathbf{x}}\right)}\right\|_{\mathcal{H}_{h_{p}}}=:K<\infty$). Then, for any $\delta\in(0,1)$ it holds that
$\left(\mathbb{Q}^{n}\otimes\Lambda^{m}\right)\Bigg{(}\left|S_{p}(\mathbb{Q})-\tilde{S}_{p}\left(\hat{\mathbb{Q}}_{n}\right)\right|\leq\frac{c_{1}}{\sqrt{n}}+\frac{c_{2}}{m}+\frac{c_{3}\sqrt{\log\frac{m}{\delta}}}{m}\sqrt{\mathcal{N}_{\mathbb{Q},h_{p}}\!\left(\frac{12K^{2}\log\frac{m}{\delta}}{m}\right)}\Bigg{)}\geq 1-\delta,$ (16)
@@
Consider univariate data ($d=1$), unnormalized target density $p(x)=e^{-x^{2}/2}$, and the Gaussian kernel $k(x,y)=e^{-\gamma(x-y)^{2}}$. Notice that, by using (4), we have
$\left\|{\xi_{p}(x)}\right\|_{\mathcal{H}_{k}^{d}}^{2}=h_{p}(x,x)=x^{2}+\mathrm{const}\stackrel{{\scriptstyle x\to\infty}}{{\to}}\infty.$ (17)
@@
In fact, a more general result holds: If one considers a bounded continuously differentiable translation-invariant kernel $k$, the induced Stein kernel is only bounded provided that the target density $p(\mathbf{x})$ has tails that are no thinner than $e^{-\sum_{i=1}^{d}|x_{i}|}$ [Hagrass et al., 2024, Remark 2], which clearly rules out Gaussian targets.
@@
We assume that the Stein feature map $h_{p}\left(\cdot,{X}\right)$ with the sampling distribution $\mathbb{Q}\in\mathcal{M}_{1}^{+}\left(\mathbb{R}^{d}\right)$ is sub-Gaussian in the sense of (2), that is, $\left\|{\left\langle{h_{p}\left(\cdot,{X}\right),u}\right\rangle_{\mathcal{H}_{h_{p}}}}\right\|_{\psi_{2}}\lesssim\left\|{\left\langle{h_{p}\left(\cdot,{X}\right),u}\right\rangle_{\mathcal{H}_{h_{p}}}}\right\|_{L_{2}(\mathbb{Q})}<\infty$ for all $u\in\mathcal{H}_{h_{p}}$.
@@
Theorem 2.
Let Assumption 1 hold, $C_{\mathbb{Q},h_{p}}\neq 0$, assume the setting of Lemma 1, $m\geq 4$ Nyström samples, and $\mathbb{P}=\mathbb{Q}$. Then, for any $\delta\in(0,1)$ it holds that
$\left(\mathbb{Q}^{n}\otimes\Lambda^{m}\right)\Bigg{(}$ $\left|S_{p}(\mathbb{Q})-\tilde{S}_{p}\left(\hat{\mathbb{Q}}_{n}\right)\right|\lesssim\frac{\sqrt{\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)}\log(6/\delta)}{n}+\sqrt{\frac{\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)\log(6/\delta)}{n}}+$ (18)
$\quad+\frac{\sqrt{\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)\log(12n/\delta)\log(12/\delta)}}{m}\sqrt{\mathcal{N}_{\mathbb{Q},h_{p}}\left(\frac{c\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)}{m}\right)}\Bigg{)}\geq 1-\delta$ (19)
when $m\gtrsim\max\left\{\left\|{C_{\mathbb{Q},h_{p}}}\right\|_{\mathrm{op}}^{-1}\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right),\log(3/\delta)\right\}$, where $c>0$ is a constant.
To interpret the consistency guarantee of Theorem 2, we consider the three terms on the r.h.s. w.r.t. the magnitude of $m$. The first two terms converge with $\mathcal{O}\left(n^{-1/2}\right)$, independent of the choice of $m$. By using the universal upper bound $\mathcal{N}_{\mathbb{Q},h_{p}}\left(\frac{c\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)}{m}\right)\lesssim m$ on the effective dimension, the last term reveals that an overall rate of $\mathcal{O}\left(n^{-1/2}\right)$ can only be achieved with further assumptions regarding the rate of decay of the effective dimension if one also requires $m=o\left(n^{2/3}\right)$ — as is necessary for a speed-up, see Remark 1(a). Indeed, the rate of decay of the effective dimension can be linked to the rate of decay of the eigenvalues of the covariance operator [Della Vecchia et al., 2021, Proposition 4, 5], which is known to frequently decay exponentially, or, at least, polynomially. In this sense, the last term acts as a balance, which takes the characteristics of the data and of the kernel into account.
@@
In the setting of Theorem 2, assume that the spectrum of the covariance operator $C_{\mathbb{Q},h_{p}}$ decays either (i) polynomially, implying that $\mathcal{N}_{\mathbb{Q},h_{p}}(\lambda)\lesssim\lambda^{-\gamma}$ for some $\gamma\in(0,1]$, or (ii) exponentially, implying that, $\mathcal{N}_{\mathbb{Q},h_{p}}(\lambda)\lesssim\log(1+\frac{c_{1}}{\lambda})$ for some $c_{1}>0$. Then it holds that
@@
assuming that the number of Nyström points satisfies (i) $m\gtrsim n^{\frac{1}{2-\gamma}}\log^{\frac{1}{2-\gamma}}(12n/\delta)\log^{\frac{1}{2-\gamma}}(12/\delta)$ in the first case, or (ii) $m\gtrsim\sqrt{n}\left(\log\left(1+\frac{c_{1}n}{c\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)}\right)\log(12n/\delta)\log(12/\delta)\right)^{1/2}$ in the second case.
@@
    Comparison of Theorem 1 and Theorem 2. The key takeaway of Theorem 2 is that the boundedness requirement on the kernel can be relaxed to a sub-Gaussian assumption for KSD estimation.
@@
    Sub-Gaussian assumption. Key to the proof of Theorem 2 is having an adequate notion of non-boundedness of the feature map. One approach—common for controlling unbounded real-valued random variables— is to impose a sub-Gaussian assumption. In Hilbert spaces, various definitions of sub-Gaussian behavior have been investigated [Talagrand, 1987, Fukuda, 1990, Antonini, 1997]; see Giorgobiani et al. [2020] for a recent survey. Classically, these require centered random variables. We note that with (6) the mean-zero requirement is satisfied for $\mathbb{P}=\mathbb{Q}$. Among the definitions of sub-Gaussianity, we carefully selected Koltchinskii and Lounici [2017, Def. 2]. [^5] The condition is also referred to as sub-Gaussian in Fukuda’s sense [Giorgobiani et al., 2020, Def. 1]. Specifically, the condition allows us to derive our key Lemma B.3, which intuitively states that norms of sub-Gaussian vectors whitened by $C_{\mathbb{Q},h_{p},\lambda}^{-1/2}$ inherit the sub-Gaussian property, opening the door for proving Theorem 2.

### v4

@@ 4.2 Nyström Bootstrap Testing
$w_{i}=\mathds{1}_{(U_{i}>0.5)}w_{i-1}-\mathds{1}_{(U_{i}\leq 0.5)}w_{i-1},$ (16)
where $U_{i}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\operatorname{Unif}(0,1)$, that is, $w_{i}$ changes sign with probability $0.5$. The test procedure is as follows.
@@
with $\mathbf{w}=\left(w_{i}\right)_{i=1}^{n}\in\mathbb{R}^{n}$ collecting the $w_{i}$-s ($i\in[n]$) defined in (16); $\mathbf{K}_{h_{p},n,m}$ $(=\mathbf{K}_{h_{p},m,n}^{\mkern-1.5mu\mathsf{T}})$ and $\mathbf{K}_{h_{p},m,m}$ are defined as in Lemma 1. The approximation is based on the fact [Williams and Seeger, 2001] that $\mathbf{K}_{h_{p},n,m}\mathbf{K}_{h_{p},m,m}^{-}\mathbf{K}_{h_{p},m,n}$ is a low-rank approximation of $\mathbf{K}_{h_{p},n,n}$, that is, $\mathbf{K}_{h_{p},n,m}\mathbf{K}_{h_{p},m,m}^{-}\mathbf{K}_{h_{p},m,n}\approx\mathbf{K}_{h_{p},n,n}$. Hence, our proposed procedure (17) approximates (15) but reduces the cost from $\mathcal{O}\left(n^{2}\right)$ to $\mathcal{O}\left(nm+m^{3}\right)$ if one computes from left to right (also refer to Remark 1(a)); in the case of $m=o\left(n^{2/3}\right)$ this guarantees an asymptotic speedup. We obtain a total cost of $\mathcal{O}\left(D\left(nm+m^{3}\right)\right)$ for obtaining the wild bootstrap samples. This acceleration allows KSD-based goodness-of-fit tests to be applied on large data sets.
@@ 4.3 Guarantees
Theorem 1 (Bounded case).
Assume the setting of Lemma 1, $C_{\mathbb{Q},h_{p}}\neq 0$, $m\geq 4$ Nyström samples, and a bounded Stein feature map ($\sup_{\mathbf{x}\in\mathbb{R}^{d}}\left\|{h_{p}\left(\cdot,{\mathbf{x}}\right)}\right\|_{\mathcal{H}_{h_{p}}}=:K<\infty$). Then, for any $\delta\in(0,1)$, it holds with $\left(\mathbb{Q}^{n}\otimes\Lambda^{m}\right)$-probability of at least $1-\delta$ that
$\left|S_{p}(\mathbb{Q})-\tilde{S}_{p}\left(\hat{\mathbb{Q}}_{n}\right)\right|\leq\frac{c_{1}}{\sqrt{n}}+\frac{c_{2}}{m}+\frac{c_{3}\sqrt{\log\frac{m}{\delta}}}{m}\sqrt{\mathcal{N}_{\mathbb{Q},h_{p}}\!\left(\frac{12K^{2}\log\frac{m}{\delta}}{m}\right)},$ (18)
@@
Consider univariate data ($d=1$), unnormalized target density $p(x)=e^{-x^{2}/2}$ (corresponding to $\mathbb{P}=\mathcal{N}(0,1)$), and (i) the RBF kernel $k(x,y)=\exp\left(-\gamma(x-y)^{2}\right)$ with $\gamma>0$, or (ii) the IMQ kernel $k(x,y)=\left(c^{2}+(x-y)^{2}\right)^{{-\beta}}$ with $\beta,c>0$. By using (6), direct calculation yields (i) $\left\|{\xi_{p}(\cdot,x)}\right\|_{\mathcal{H}_{k}}^{2}=x^{2}+2\gamma\stackrel{{\scriptstyle x\to\infty}}{{\to}}\infty$ in the first, and (ii) $\left\|{\xi_{p}(\cdot,x)}\right\|_{\mathcal{H}_{k}}^{2}=x^{2}c^{2\beta}-2\beta c^{2(\beta-1)}\stackrel{{\scriptstyle x\to\infty}}{{\to}}\infty$ in the second case.
@@
In fact, a more general result holds: If one considers a bounded continuously differentiable translation-invariant kernel $k$, the induced Stein kernel is only bounded provided that the target density $p(\mathbf{x})$ has tails that are no thinner than $e^{-\sum_{i=1}^{d}|x_{i}|}$ [Hagrass et al., 2025, Remark 2], which clearly rules out Gaussian targets.
@@
The centered Stein feature map $\bar{h}_{p}\left(\cdot,{X}\right)=h_{p}\left(\cdot,{X}\right)-\mathbb{E}_{X\sim\mathbb{Q}}h_{p}\left(\cdot,{X}\right)$ with the sampling distribution $\mathbb{Q}\in\mathcal{M}_{1}^{+}\left(\mathbb{R}^{d}\right)$ is sub-Gaussian in the sense of (2), that is,
$\left\|{\left\langle{\bar{h}_{p}\left(\cdot,{X}\right),u}\right\rangle_{\mathcal{H}_{h_{p}}}}\right\|_{\psi_{2}}\lesssim\left\|{\left\langle{\bar{h}_{p}\left(\cdot,{X}\right),u}\right\rangle_{\mathcal{H}_{h_{p}}}}\right\|_{L_{2}(\mathbb{Q})}<\infty$ (19)
holds for all $u\in\mathcal{H}_{h_{p}}$, with a $u$-independent absolute constant in $\lesssim$.
Example 2 (Applicability of Assumption 1).
In the simple case $d=1$, $k(x,y)=xy$ ($\mathcal{H}_{k}=\mathbb{R}$), and target measure $\mathbb{P}=\mathcal{N}(0,1)$, Assumption 1 is satisfied, for instance, for $\mathbb{Q}=\operatorname{Unif}\left(-\sqrt{3},\sqrt{3}\right)$. The details are as follows. From (4), $\xi_{p}(\cdot,x)=h_{p}(\cdot,x)=1-x^{2}$ ($x\in\mathbb{R}$). We note that $\mathbb{E}_{X\sim\mathbb{Q}}h_{p}(\cdot,X)=0$ implies that $\bar{h}_{p}(\cdot,x)=h_{p}(\cdot,x)$ and thus we obtain $\left\|{\left\langle{h_{p}(\cdot,X),u}\right\rangle_{\mathbb{R}}}\right\|_{\psi_{2}}=|u|\left\|{1-X^{2}}\right\|_{\psi_{2}}\stackrel{{\scriptstyle(a)}}{{\leq}}|u|c_{1}\stackrel{{\scriptstyle(b)}}{{=}}|u|c_{1}c_{2}\left\|{1-X^{2}}\right\|_{L_{2}(\mathbb{Q})}\lesssim\left\|{\left\langle{h_{p}(\cdot,X),u}\right\rangle_{\mathbb{R}}}\right\|_{L_{2}(\mathbb{Q})}$. The boundedness of $X$ implies the sub-Gaussianity (in the real-valued sense) of $1-X^{2}$ in (a); hence, $\left\|{1-X^{2}}\right\|_{\psi_{2}}\leq c_{1}$. In (b), we let $c_{2}=\left\|{1-X^{2}}\right\|_{L_{2}(\mathbb{Q})}^{-1}$.
@@
Theorem 2 (Consistency of Nyström-KSD).
Let Assumption 1 hold, $C_{\mathbb{Q},\bar{h}_{p}}\neq 0$, and assume the setting of Lemma 1. Then, for any $\delta\in(0,1)$ with $\left(\mathbb{Q}^{n}\otimes\Lambda^{m}\right)$-probability of at least $1-\delta$ it holds that
$\left|S_{p}(\mathbb{Q})-\tilde{S}_{p}\left(\hat{\mathbb{Q}}_{n}\right)\right|$ $\lesssim\frac{\sqrt{\operatorname{tr}\left(C_{\mathbb{Q},\bar{h}_{p}}\right)}\log(6/\delta)}{n}+\sqrt{\frac{\operatorname{tr}\left(C_{\mathbb{Q},\bar{h}_{p}}\right)\log(6/\delta)}{n}}$ (20)
$\quad+\frac{\sqrt{\operatorname{tr}\left(C_{\mathbb{Q},\bar{h}_{p}}\right)\log(12n/\delta)\log(12/\delta)}}{m}\sqrt{\mathcal{N}_{\mathbb{Q},\bar{h}_{p}}\left(\frac{c\operatorname{tr}\left(C_{\mathbb{Q},\bar{h}_{p}}\right)}{m}\right)}$ (21)
when $m\gtrsim\max\left\{\left\|{C_{\mathbb{Q},\bar{h}_{p}}}\right\|_{\mathrm{op}}^{-1}\operatorname{tr}\left(C_{\mathbb{Q},\bar{h}_{p}}\right),\log(12/\delta)\right\}$, where $c>1$ is a constant.
To interpret the consistency guarantee of Theorem 2, we consider the three terms on the r.h.s. w.r.t. the magnitude of $m$. The first two terms converge with $\mathcal{O}\!\left(n^{-1/2}\right)$, independent of the choice of $m$. By using the universal upper bound $\mathcal{N}_{\mathbb{Q},\bar{h}_{p}}\left(\frac{c\operatorname{tr}\left(C_{\mathbb{Q},\bar{h}_{p}}\right)}{m}\right)\lesssim m$ on the effective dimension, the last term reveals that an overall rate of $\mathcal{O}\!\left(n^{-1/2}\right)$ can only be achieved with further assumptions regarding the rate of decay of the effective dimension if one also requires $m=o\left(n^{2/3}\right)$ — as is necessary for a speed-up, see Remark 1(a). Indeed, the rate of decay of the effective dimension can be linked to the rate of decay of the eigenvalues of the covariance operator [Della Vecchia et al., 2021, Proposition 4, 5], which is known to frequently decay exponentially, or, at least, polynomially. In this sense, the last term acts as a balance, which takes the characteristics of the data and of the kernel into account.
@@
In the setting of Theorem 2, assume that the spectrum of the covariance operator $C_{\mathbb{Q},\bar{h}_{p}}$ decays either (i) polynomially, implying that $\mathcal{N}_{\mathbb{Q},\bar{h}_{p}}(\lambda)\lesssim\lambda^{-\gamma}$ for some $\gamma\in(0,1]$, or (ii) exponentially, implying that, $\mathcal{N}_{\mathbb{Q},\bar{h}_{p}}(\lambda)\lesssim\log(1+\frac{c_{1}}{\lambda})$ for some $c_{1}>0$. Then it holds that
@@
assuming that the number of Nyström points satisfies (i) $m\gtrsim n^{\frac{1}{2-\gamma}}\log^{\frac{1}{2-\gamma}}(12n/\delta)\log^{\frac{1}{2-\gamma}}(12/\delta)$ in the first case, or (ii) $m\gtrsim\sqrt{n}\left(\log\left(1+\frac{c_{1}n}{c\operatorname{tr}\left(C_{\mathbb{Q},\bar{h}_{p}}\right)}\right)\log(12n/\delta)\log(12/\delta)\right)^{1/2}$ in the second case.
To interpret these rates—see Remark 3(d)—, we obtain the (matching) $\sqrt{n}$-consistency of the quadratic time estimator (10) in our following result.
Theorem 3 (Consistency of KSD).
Assume that $\left\|{\left\|{h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{\psi_{2}}<\infty$ and define ^$\hat{\mathbb{Q}}_{n}=\{X_{1},\ldots,X_{n}\}$, where $\{X_{i}\}_{i\in[n]}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathbb{Q}$. Then it holds that
$\left|S_{p}(\mathbb{Q})-S_{p}\left(\hat{\mathbb{Q}}_{n}\right)\right|=\mathcal{O}_{\mathbb{Q}^{n}}\left(\frac{1}{\sqrt{n}}\right).$ (23)
The following example illustrates that, in some cases, the assumption $\left\|{\left\|{h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{\psi_{2}}<\infty$ can be verified analytically.
Example 3 (Assumption $\left\|{\left\|{h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{\psi_{2}}<\infty$).
Assume that $d=1$, $k=\exp\left(-\gamma(x-y)^{2}\right)$ ($\gamma>0$), target measure $\mathbb{P}=\mathcal{N}(0,1)$, and samples $X,X_{1},\ldots,X_{n}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathbb{Q}$ with $\left\|{X}\right\|_{\psi_{2}}<\infty$. Then
$\left\|{\left\|{h_{p}(\cdot,X)}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{\psi_{2}}^{2}$ =$\stackrel{{\scriptstyle(a)}}{{=}}\left\|{h_{p}(X,X)}\right\|_{\psi_{1}}\stackrel{{\scriptstyle(b)}}{{=}}\left\|{X^{2}+2\gamma}\right\|_{\psi_{1}}$
$\stackrel{{\scriptstyle(c)}}{{\leq}}\left\|{X^{2}}\right\|_{\psi_{1}}+\left\|{2\gamma}\right\|_{\psi_{2}}\stackrel{{\scriptstyle(d)}}{{=}}\left\|{X}\right\|_{\psi_{2}}+\frac{2\gamma}{\sqrt{\log 2}}<\infty,$
with the following details. Lemma C.2(iv) and the reproducing property yield (a). (b) follows from the explicit form of $h_{p}$ given in Example 1(i). The triangle inequality gives (c) and (d) follows from the definition of the $\psi_{2}$-norm using that $2\gamma$ is non-random.
In this setting, similar computations using Example 1(ii) show that the assumption is also satisfied with the IMQ kernel.
A few remarks are in order.
@@
    Comparison of Theorem 1 and Theorem 2. Recall that both theorems target precisely the same estimators, Chatalic et al. [2022, (5)] and (14), respectively. We note that in the finite-dimensional case, every bounded random variable is also sub-Gaussian. This property does not carry over to sub-Gaussianity in the infinite-dimensional case; see the remark after Della Vecchia et al. [2021, Definition 1]. In this sense, the assumptions of both statements are not directly comparable. Still, the takeaway of both results—with these different sets of conditions—is the same.
@@
    Sub-Gaussian assumption. Key to the proof of Theorem 2 is having an adequate notion of non-boundedness of the feature map. One approach—common for controlling unbounded real-valued random variables— is to impose a sub-Gaussian assumption. In Hilbert spaces, various definitions of sub-Gaussian behavior have been investigated [Talagrand, 1987, Fukuda, 1990, Antonini, 1997]; see Giorgobiani et al. [2020] for a recent survey. Among the definitions of sub-Gaussianity, we carefully selected Koltchinskii and Lounici [2017, Def. 2]. [^6] The condition is also referred to as sub-Gaussian in Fukuda’s sense [Giorgobiani et al., 2020, Def. 1]. Specifically, this assumption allows us to derive our key Lemma B.1 and Lemma B.3. The former is similar to Rudi et al. [2015, Lemma 6], which is typically employed for Nyström analysis in the bounded case [Chatalic et al., 2022, Sterge and Sriperumbudur, 2022, Kalinke and Szabó, 2023], but our result applies to the sub-Gaussian setting. The main technical challenge we resolve is transforming our setting to a form in which existing concentration results can be leveraged. Especially the case of $\mathbb{P}\neq\mathbb{Q}$ requires special care, which we tackle by systematically using the centered covariance operator $C_{\mathbb{Q},\bar{h}_{p}}$; we refer to the respective proof for details. [^7] We note that an analysis of the centered setting is also challenging in the bounded case; for instance, Sterge and Sriperumbudur [2022] tackle the resulting difficulties (in case of kernel PCA) with U-statistics, of which our method is independent. The latter, Lemma B.3, intuitively states that norms of sub-Gaussian vectors whitened by $C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}$ inherit the sub-Gaussian property. Together, these lemmas open the door to proving Theorem 2.
  4. (d)
    Comparison of Theorem 2 and Theorem 3. With the weaker RKHS norm condition $\left\|{\left\|{h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{\psi_{2}}<\infty$ (implied by Assumption 1, see Lemma B.3), Theorem 3 shows that the quadratic time estimator (10) converges with rate $\mathcal{O}\!\left(n^{-1/2}\right)$. Our Nyström result, Theorem 2 with Corollary 1, shows that a matching rate can be achieved (given an appropriate decay of the effective dimension) with $m=\tilde{\Theta}\!\left(\sqrt{n}\right)$; this choice of $m$ satisfies $m=o\left(n^{2/3}\right)$ and thus implies an asymptotic speedup by (a).
  5. (e)
    General KSD framework. We note that our results also hold in the general KSD framework [Hagrass et al., 2025] but we present them on $\mathbb{R}^{d}$, which one arguably most frequently encounters in practice, to simplify exposition.